An asset has a value of $1,000,000. In an attack, it is expected to lose 60 percent of its value. Countermeasure X will cut the loss by two-thirds. Countermeasure Y will cut the loss by half. Both countermeasures will cost $20,000 per year. An attack is likely to be successful once in 10 years. Both countermeasures can cut the occurrence rate in half. Do an analysis of these countermeasures and then give your recommendation.
Answer #1

Original Loss Due to attack = 60%*1000000=$600000

In counter measure X: the attack will happen once is 20 years and loss will be 1/3*600000=$200000

Total cost for 20 years = 20000*20=$2,40,000

Hence total cost =240000+200000=$440000

Hence Net Benefit = 600000-440000=$160000

In counter measure Y: the attack will happen once is 20 years and loss will be 1/2*600000=$300000

Total cost for 20 years = 20000*20=$2,40,000

Hence total cost =240000+300000=$540000

Hence Net Benefit = 600000-540000=$60000

Hence the Net Benefit is more in case of Countermeasure X hence X is recommended